4 After the accounts have been adjusted at July 31, the end of the fiscal year, the ollowing 390.450 Sales The cost of sales 280,150 Wages Expense 70,000 20,000 Miscellaneous Expense &8200 Josmalize the entries required to close the accousts Eipmen squired at the beginning of the year at a oost of $125,000 hs as esimuted residual value of $5,000 and an estimated useful life of 5 years. Die yars depreciation (23 pereent)
Answer #1

[As HOMEWORKLIB’s policy and guideline, the first-question with all parts is answered below:]

Q3)

a.

SL depreciation = (Cost – Residual value) / Life years

                           = ($125,000 - $5,000) / 5

                           = $120,000 / 5

                           = $24,000 (Answer)

b.

DB rate = SL rate × 2

            = [(1/Life years) × 100] × 2

            = [(1/5) × 100] × 2

            = 20% × 2

            = 40% (Answer)

c.

DB depreciations (every year):

Year 1 = Cost × DB rate = $125,000 × 40% = $50,000 (Answer)

Year 2 = (Cost – 50,000) × DB rate = ($125,000 – 50,000) × 40% = $75,000 × 40% = $30,000 (Answer)

Year 3 = (Cost – 50,000 – 30,000) × DB rate = ($125,000 – 80,000) × 40% = $45,000 × 40% = $18,000 (Answer)

Year 4 = (Cost – 50,000 – 30,000 – 18,000) × DB rate = ($125,000 – 98,000) × 40% = $27,000 × 40% = $10,800 (Answer)

Year 5 = (Cost – 50,000 – 30,000 – 18,000 – 10,800) × DB rate = ($125,000 – 108,800) × 40% = $16,200 × 40% = $6,480. But the residual value of $5,000 is to be maintained in the last year. Therefore, the depreciation is ($16,200 - $5,000 =) $11,200 (Answer)

d.

Journal entry

Date

Account titles & explanation

P.ref

Debit $

Credit $

End of year 1

Depreciation expense

50,000

      Accumulated depreciation

50,000

To record charging depreciation of equipment at the year-end.